OpenStudy (anonymous):

Raising a quantity to a negative exponent will produce __________. A. a decimal B. a negative number C. the reciprocal of the positive power D. the additive inverse of the quantity

OpenStudy (ipwnbunnies):

It'll be the reciprocal. For example: \[x^{-a} = \frac{1}{x^a}\] \[2^{-2} = \frac{1}{2^{2}} = \frac{1}{4}\]

OpenStudy (anonymous):

Thanks!